# Expedited thermalization dynamics in incommensurate systems

### v1

@@ IV Conclusion and outlook
In this work, we have systematically investigated the thermalization dynamics of quantum systems embedded in incommensurate potentials and coupled to Markovian reservoirs, uncovering a robust inverse QME where initially colder and more localized states thermalize faster to infinite temperature than their warmer and extended counterparts. Through a combination of analytical and numerical approaches, we have demonstrated that this counterintuitive behavior stems from the reduced overlap between localized states and the slowest-decaying Liouvillian modes, providing a clear physical mechanism for the accelerated thermalization. The effect persists across parameter variations and system sizes, indicating its fundamental nature in dissipative quantum systems with localization properties.
@@ Appendix A Proof of maximally mixed steady-state under hermitian dissipation operators
In this appendix, we prove the conclusion that a steady-state is maximally mixed state when dissipation operators are hermitian. For simplicity, we assume the dissipation operator is local dephasing, namely, $O_{j}=n_{j}$. Defining the matrix element $\rho_{i,j}(t)=\langle{i}|\rho(t)|{j}\rangle$ in the local basis, where $|j\rangle=\hat{c}_{j}^{\dagger}|0\rangle$ is the single-particle state at the lattice site $j$, we can obtain the time evolution equations from the Lindblad master equation
$\frac{\mathrm{d}}{\mathrm{~{}d}t}\rho_{i,j}=$ $-iJ\left(\rho_{i,j+1}+\rho_{i,j-1}-\rho_{i+1,j}-\rho_{i-1,j}\right)$ (11)
@@
where $\delta_{i,j}$ denotes the (Kronecker) delta function. We can prove that the steady-state solution $\hat{\rho}_{\rm ss}$ to above equation which corresponds to $\frac{\mathrm{d}}{\mathrm{~{}d}t}\rho_{\it{i,j}}=0$ for any matrix element, is the maximally-mixed state, namely the steady-state density matrix $\rho_{\rm ss}$ is proportional to the identity matrix

### v2

@@ IV Conclusion and outlook
In this work, we have systematically investigated the thermalization dynamics of quantum systems embedded in incommensurate potentials and coupled to Markovian reservoirs, uncovering a robust inverse QME phenomenon where initially colder and more localized states thermalize faster to infinite temperature than their warmer and extended counterparts. Through a combination of analytical and numerical approaches, we have demonstrated that this counterintuitive behavior stems from the reduced overlap between localized states and the slowest-decaying Liouvillian modes, providing a clear physical mechanism for the accelerated thermalization. The effect persists across parameter variations and system sizes, indicating its fundamental nature in dissipative quantum systems with localization properties.
@@ Appendix A Proof of maximally mixed steady-state under hermitian dissipation operators
In this appendix, we prove the conclusion that a steady-state is maximally mixed state when dissipation operators are hermitian. For simplicity, we assume the dissipation operator is local dephasing, namely, $O_{j}=n_{j}$ and consider a generic tight-binding model. Defining the matrix element $\rho_{i,j}(t)=\langle{i}|\rho(t)|{j}\rangle$ in the local basis, where $|j\rangle=\hat{c}_{j}^{\dagger}|0\rangle$ is the single-particle state at the lattice site $j$, we can obtain the time evolution equations from the Lindblad master equation
$\frac{\mathrm{d}}{\mathrm{~{}d}t}\rho_{i,j}=$ $-iJ\left(\rho_{i,j+1}+\rho_{i,j-1}-\rho_{i+1,j}-\rho_{i-1,j}\right)$
@@
where $\epsilon_{\it j}$ ($\epsilon_{\it i}$) is the onsite potential at the site $j$ ($i$), $J$ is the hopping strength, and $\delta_{i,j}$ denotes the (Kronecker) delta function. We can prove that the steady-state solution $\rho_{\rm ss}$ to above equation which corresponds to $\frac{\mathrm{d}}{\mathrm{~{}d}t}\rho_{\it{i,j}}=0$ for any matrix element, is the maximally-mixed state, namely the steady-state density matrix $\rho_{\rm ss}$ is proportional to the identity matrix